The sum of the ages of Rita and Lara is 92. In 6 years, the difference of their ages will be 24. How old is each now?
Answer by Cromlix(4381)

Hi there,
The sum of the ages of Rita and Lara is 92.
R + L = 92
R = 92 - L
In 6 years, the difference of their ages will be 24.
(R + 6) - (L + 6) = 24
Remove brackets
R + 6 - L - 6 = 24
Replace R with 92 - L (Top equation)
92 - L + 6 - L - 6 = 24
Collect like terms
-2L = 24 - 92
-2L = - 68
L = 34
Lara = 34 yrs old
Rita = 58 yrs old
58 + 6 - 34 + 6
64 - 40 = 24
Hope this helps :-)